OpenStudy (anonymous):

will give medal Presley is organizing the school book sale. Each stall will have 40 books on the shelf and 2 boxes of books with b books inside. Presley also has 200 books in inventory to refill the shelves as books sell. This situation can be written as the following expression: 40s + (2b)s + 200 How many books should Presley have ready if she plans to open 8 stalls and have 12 books in each box? (2 points) 520 712 820 1,622

OpenStudy (anonymous):

T=8(40+2(12))+200 T=8(40+24)+200 T=8(64)+200 T=512+200=712 She needs to have 712 books on hand to open 8 stalls.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

NP ^^